OpenStudy (anonymous):

Find the LCM of 21x^2 + 105x and 3x^2 + 24x + 45

OpenStudy (anonymous):

6x^2 + 30x can factor to 6x(x+5) 3x^2 + 18x + 15 can factor to (x+5)(3x+3) so the least common multiple is (x+5)

OpenStudy (asnaseer):

@thechocoluver445 you have found the greatest common divisor (GCD). the LCM is different.

OpenStudy (anonymous):

well 3x^2 + 18x + 15 = 3(x+1)(x+5) so the lcm is 3. my bad.

OpenStudy (asnaseer):

This is still wrong I'm afraid. @byndbtch if you want to learn about this there is a good explanation here: http://www.purplemath.com/modules/lcm_gcf.htm

OpenStudy (anonymous):

oh, would the lcm be 3(x+5)? sorry, i have done this in a while..

OpenStudy (asnaseer):

you correctly found that: 6x^2 + 30x can factor to 6x(x+5) = 2 * 3 * x * (x+5) 3x^2 + 18x + 15 can factor to (x+5)(3x+3) = 3 * (x+5) * (x+1) now if you compare the two, you need to find the product of all "unique" factors of both numbers. in this case the LCM = 2*3*x*(x+5)*(x+1) = 6x(x+5)(x+1)

OpenStudy (anonymous):

how about this one?? 35y^2 + 175y and 7y^2 + 42y + 35

OpenStudy (asnaseer):

Use the same principals - that link I gave above explains LCM in a lot of detail.

OpenStudy (anonymous):

ty but it is not sinking in. I appreciate your help.

OpenStudy (asnaseer):

it might help if you at least try to solve it. just post the steps that you /think/ you need to take and I will guide you if necessary to the correct solution.

OpenStudy (asnaseer):

e.g. the first step is to factorise each polynomial

OpenStudy (anonymous):

21y^2 + 84y = 21y(y +21)

OpenStudy (anonymous):

thats not right

OpenStudy (asnaseer):

its /almost/ right, HINT: 84 = 21 * 4

OpenStudy (anonymous):

21y(y + 4) ??

OpenStudy (asnaseer):

perfect! now factorise the other polynomial as well please

OpenStudy (anonymous):

21y(y +4)(y +1)

OpenStudy (asnaseer):

BTW: What are the two polynomials you are working with here?

OpenStudy (asnaseer):

I'm asking because I noticed that you have changed the original polynomials that you posted in this question.

OpenStudy (anonymous):

21y^2 + 84y and then 7y^2 + 42y +56

OpenStudy (asnaseer):

ok, so you factorised the first one correctly: 21y^2 + 84y = 21y(y + 4) but the second factorisation does not look correct for 7y^2 + 42y +56

OpenStudy (asnaseer):

do you want to try again with that one?

OpenStudy (anonymous):

yeah sorry about that, i was rushing and kept confusing gcf with lcf with lcm...sorry :/

OpenStudy (asnaseer):

@thechocoluver445 np :)

OpenStudy (anonymous):

acco\rding to my question, the answer is 21y(y+4)(y +2) which I have no idea where the 2 came from????

OpenStudy (asnaseer):

ignore the answer for now. lets look at the second polynomial which is: 7y^2 + 42y +56 notice all the terms can be divided by 7

OpenStudy (anonymous):

right

OpenStudy (asnaseer):

so 7y^2 + 42y +56 = 7(y^2+6y+8) can you factorise y^2+6y+8?

OpenStudy (anonymous):

not really....as I do notunderstand where the 6and 8 came from

OpenStudy (asnaseer):

7*6 = 42 7*8 = 56 so: 7(y^2+6y+8) = 7y^2 + (7*6)y + (7*8) = 7y^2 + 42y + 56

OpenStudy (asnaseer):

does it make any sense now?

OpenStudy (anonymous):

no it really doesnt but I do appreciate the effort....Im going to take a break now....thanks again.

OpenStudy (asnaseer):

np - let me know if you need more help at a later date and I'll be more than happy to help :)